OpenStudy (anonymous):

A problem states: "There are 5 more girls than boys in a class. There are 27 students in the class in all. How many boys are there in the class?" What are the unknowns in this problem?

OpenStudy (midhun.madhu1987):

Let number of boys be x. So, number of girls = x + 5 Total students = 27 So, x + (x+5) = 27 Find x.

OpenStudy (anonymous):

what is x

OpenStudy (anonymous):

im finding the unknowns of the problem

OpenStudy (anonymous):

no it will just be x+5=27

OpenStudy (anonymous):

the answer choices are the number of students in the class the number of boys in the class the number of girls in the class the number of boys and the number of girls in the class

OpenStudy (anonymous):

can u explain why will it be 2x

OpenStudy (midhun.madhu1987):

Is the number of boys known??

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

tell me if it a , b, c or d

OpenStudy (midhun.madhu1987):

Is the number of boys known in the question??